40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 19 23:50:44 2013, Jenne, Update, LSC, PRCL locking again 
    Reply  Thu Jun 20 01:13:18 2013, Jenne, Update, LSC, PRCL locking again - no ASC success 
       Reply  Wed Jun 26 04:38:02 2013, Jenne, Update, LSC, PRCL locking again - ASC success PRMI_ASC_yawOnly_powerSpectra_25June2013.pdfPRMI_ASC_yawOnly_25June2013_mag.pdfPRMI_ASC_yawOnly_25June2013_phase.pdfPRMI_ASC_yawOnly_25June2013_coherence.pdfSensMatMeas_26June2013.png
          Reply  Thu Jun 27 17:09:41 2013, Jenne, Update, LSC, PRCL locking again - POP PIT work MeasuredVsModeledASCyaw.png
             Reply  Fri Jun 28 02:12:41 2013, Jenne, Update, LSC, PRCL locking again - ASC work PRMI_ASC_powerSpectra_27June2013.pdfScreenshot-Untitled_Window.png
Message ID: 8780     Entry time: Fri Jun 28 02:12:41 2013     In reply to: 8767
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRCL locking again - ASC work 

Rana had the epiphany that I didn't have any antiwhitening for my POP QPD.  Ooops. 

We looked at the schematic for the Pentek Generic board (pdf), and saw that it has a Zero @ 15Hz, and Poles @ 150Hz and 1500Hz, times 2 stages.  We determined from the TF that I posted that probably both stages are engaged, so I made an antiwhitening filter consisting of the inverse (so, 2 poles at 15Hz, 2 zeros at 150Hz and 2 zeros at 1500Hz).  [Rana points out that for this low frequency system we may not want to include the 1500Hz compensation, since it is probably just enhancing ADC noise].  The ASC system worked really well, really easily, after that.

Another note though, the AA stage of the Pentek Generic boards have 4 poles at 800Hz, which are not compensated.

Rana also added a 60Hz comb to the filter bank with the AntiWhitening, since the QPD has an unfortunately large amount of 60Hz noise.  Also, the 60Hz lowpass in the ASC loop was engaged for both pitch and yaw.

Rana, Lisa and Manasa also found that the ASC system was *more* stable with the PRM oplev ON. 

So, the ASC locking situation is:

PRM oplev loops on.

AS-POP_QPD_[PIT/YAW] filter banks with FM1, FM6 on.

ASC-PRCL_[PIT/YAW] filter banks with FM1, FM5, FM6 and FM9 on.



(No triggering yet).

The ASC Up and Down scripts (which are called from the buttons on the ASC screen) have all of these gain settings, although they assume for now that all the filters are already on.

Here's a screenshot of the power spectra showing the angular motion suppression. The PDF is attached so you can zoom in and see some details.  The dashed lines are the "PRMI locked, ASC off" case, and the solid lines are the "PRMI locked, ASC on" case.  You can see that according to the QPD, we do an excellent job suppressing both the pitch and yaw motion (although better for yaw), but there isn't a huge effect on POPDC or POP110I.  While we could probably do better if we had a 2 QPD system with the QPDs at differet gouy phases, this seems to be good enough that we can keep the PRMI locked ~indefinitely. 


I would like to compile the ASC model, so that I can implement triggering.  For tonight, we did not have the ASC engaged during our PRMI+Xarm tests (see Manasa's elog), but I think it'll make things a little easier if we can get the ASC going automatically.

Attachment 1: PRMI_ASC_powerSpectra_27June2013.pdf  110 kB  | Hide | Hide all
ELOG V3.1.3-